How many liters of a 40% alcohol solution must be mixed with a 60 L of a 80% solution to get a 50% solution?
Answer by stanbon(75887)

How many liters of a 40% alcohol solution must be mixed with a 60 L of a 80% solution to get a 50% solution?
---------------------
Equation:
alco + alco = alco
0.80*60 + 0.40x = 0.50(60+x)
---
Multiply by 100 and solve for "x":
80*60 + 40x = 50*60 + 50x
-10x = -30*60
x = (1/3)*60
x = 20L (amt. of 40% solution needed)
=========================================
Cheers,
Stan H.
==============